gelphielvr:

(algebra 1) not looking for the answer. Does anyone know if this question is asking for the answer or for the equation to find the answer? (Q in the replies)

gelphielvr:

1 attachment
Twaylor:

500x2%

Twaylor:

...

Twaylor:

thats litterally it

gelphielvr:

yeah but are they asking for the answer or for the equation?

Twaylor:

equasion

gelphielvr:

@twaylor wrote:
equasion
okay thank you

dude:

I'll add a note here since the question hadn't been fully answered. The general equation of the exponential function here is \(y=a\times (1+r)^x\) where "a" is the starting point and "r" is the rate at which it increases(or decreases) The question states "company currently has 500 employees." and "number of employees is expected to grow at a rate of 2% each year" - "currently" we can understand as the starting point. In this case, we can replace the a in the equation I gave earlier with this. - Increases by 2% per year is the rate, we can add this to the r section of the equation I mentioned earlier We'll have: \(y=500(1+2\%)^x\). Although this could be an answer, you want to get rid of percentages when you deal with these questions. 2% = 0.02 (If you need a refresher as to why this is true, feel free to ask) We end up with \(\color{green}{y=500(1.02)^x}\)